OpenStudy (anonymous):

medal and fan if helped! algebra Point S is the midpoint of RT.The coordinates of points R and S are (-1,-3) and (5,-5), respectively.What are the coordintates of point T? A.(7,1) b.(2,-4) c.(7,-11) d(11,-7)

OpenStudy (michele_laino):

please you have to apply these formula: \[x _{S}=\frac{ x _{1}+x _{2} }{ 2 },y _{S}=\frac{ y _{1} +y _{2}}{ 2 }\] where x_S,y_S are the midpoint coordinates, and (x_1,y_1) (x_2,y_2), are your points

OpenStudy (anonymous):

whats Xs and Ys?

OpenStudy (anonymous):

im not looking for midpoint ,thats a,lready given @Michele_Laino

OpenStudy (michele_laino):

x_S and yS are the coordinate of your midpoint S, namely: \[S=(x _{S},y _{S})\]

OpenStudy (anonymous):

The mid point is (5,-5) but i need to find T

OpenStudy (michele_laino):

\[(x _{1},y _{1})=(-1,-3)\]

OpenStudy (michele_laino):

and: \[T=(x _{2},y _{2}) \]

OpenStudy (anonymous):

What formula do i use in order to Find T?

OpenStudy (michele_laino):

please you ahve to substitute your numerical data into my formulas, namely: \[5=\frac{ -1+x }{ 2 },\] \[-5=\frac{ -3+y }{ 2 }\] plese solve that equation for x and y

OpenStudy (anonymous):

x=6? y=-2?

OpenStudy (michele_laino):

no, \[x=2*5+1=11\]

OpenStudy (anonymous):

so y= -11?

OpenStudy (michele_laino):

and : \[y=(-5)*2+3=-10+3=-7\]

OpenStudy (michele_laino):

so, your option is D

OpenStudy (anonymous):

oh i get it Thank you so much!

OpenStudy (michele_laino):

Thank you!